OpenStudy (anonymous):

please help me with this: The resistivity of a copper rod 50ft long and 0.25 in. in diameter is 1.76 x 10 ^-6 ohm cm at 20 Celsius. What is its resistance at -20 Celsius?

OpenStudy (anonymous):

The formula is: R2 = R1 ( 1 + a ( t2 - t1 ) ) R1 = 50 (Resistance no. 1) [Ohm] t1 = 20 (Temperature no. 1) [Celsius] t2 = 80 (Temperature no. 2) [Celsius] a = 0.004 (Temperature coefficient) For more info on that one check http://www.allaboutcircuits.com/vol_1/chpt_12/6.html The resistance, R, of a length of wire is described by the expression: R = ρL/A ρ = resistivity of the material composing the wire, L = length of the wire, and A = area of the conducting cross section of the wire. I guess this will do. Have fun.